OpenStudy (anonymous):

please help! 4-(6)/-5= a. 2/5 b. -2/5 c. 2 d. -2

OpenStudy (anonymous):

none of those are right

OpenStudy (anonymous):

i got 5.2

OpenStudy (anonymous):

Really my calculator gave me like 4.8

OpenStudy (misty1212):

\[4-6=-2\\ \frac{-2}{-5}=\frac{2}{5}\]

OpenStudy (anonymous):

Misty Theres no way thats right how did u get that

OpenStudy (anonymous):

@misty1212

OpenStudy (anonymous):

Im intrigued xD

OpenStudy (anonymous):

Why

OpenStudy (anonymous):

i put it in my Cal... and i got 5.2

OpenStudy (misty1212):

\[\frac{4-6}{-5}=\frac{-2}{-5}=\frac{2}{5}\] i did it by arithmetic

OpenStudy (anonymous):

Well stupid Calculator need to get smarter

OpenStudy (anonymous):

She Right

OpenStudy (anonymous):

yeaah thanks xd

OpenStudy (anonymous):

4-6/-5 Take the negative outside = -1(4-6/5) = -2/5 Multiply it by the negative 1 = (- (2/5) ) = 2/5